OpenStudy (anonymous):

I would love some help please...At a private school in addition to tuition, students are charged for lab fees and supplies as shown: Kgtn = $230 1st = $271 2nd= $300 4th = $340 6th = $369 7th = $385 Let f(n) be the per student charge (in dollars) for lab fees and supplies for nth grade, kindergarten is shown by n=0. A model for this situation is f(n)=228 4√n+1 Estimate for what grade the per student charge for lab fees and supplies is $394.

OpenStudy (dumbcow):

is that sqrt(n) or sqrt(n+1)

OpenStudy (anonymous):

n+1

OpenStudy (anonymous):

sorry should have put (n+1)

OpenStudy (dumbcow):

whats in between the 8 and 4? could you rewrite f(n)

OpenStudy (anonymous):

= 228 4√n+1) There is a space btwn 228 then a small raised 4 before the radical

OpenStudy (dumbcow):

oh ok f(n) = 394, solve for n \[\sqrt[4]{n+1} = \frac{394}{228}\] \[n+1 = (1.728)^{4}\] \[n=7.92\]

OpenStudy (dumbcow):

so in other words, 8th grade :)

OpenStudy (anonymous):

but wait i get 8.916 for (1.728)^4

OpenStudy (dumbcow):

yes but that equals n+1 subtract 1 on both sides...

OpenStudy (anonymous):

got it...thank you si much

OpenStudy (anonymous):

so much